OpenStudy (yash2651995):

what is: 0^{0} ie zero to the power zero and why ? what is: 0^{0} ie zero to the power zero and why ? @Mathematics

OpenStudy (anonymous):

1

OpenStudy (anonymous):

according to some Calculus textbooks, 0^0 is an ``indeterminate form''. When evaluating a limit of the form 0^0, then you need to know that limits of that form are called ``indeterminate forms'', and that you need to use a special technique such as L'Hopital's rule to evaluate them. Otherwise, 0^0 = 1 seems to be the most useful choice for 0^0.

OpenStudy (anonymous):

It can be sometimes 1 or indeterminate, it really depends on the person defining it. See the Wikipedia article for discussion.

OpenStudy (asnaseer):

\[x ^{a}*x ^{b}=x ^{a+b}\]so if a is 1 and b is -1 we get\[x ^{1}*x ^{-1}=x ^{1-1}\]\[x * 1/x = x ^{0}\]\[x/x=x ^{0}\]\[1=x ^{0}\]

OpenStudy (asnaseer):

set x=0 in the above and you get\[0 ^{0}=1\]

OpenStudy (yash2651995):

a^0=1......(1) 0^b=0......(2) if a=0, b=0 which equation should I choose? (1) or (2)

OpenStudy (asnaseer):

from my example above\[x ^{0}=x/x\]so you can take limits to get\[0 ^{0}=\lim_{x \rightarrow 0}(x/x)=1\]

OpenStudy (asnaseer):

yash2651995 your equation above where you said 0^b = 0 is ONLY true for b > 0. e.g. if b=-1 you would get\[0 ^{-1}=1/0=\infty\]

OpenStudy (asnaseer):

so for positive n:\[0 ^{n}=0\]\[0 ^{0}=1\]\[0 ^{-n}=\infty\]

OpenStudy (zarkon):

undefined